# A framework for LISA population inference

### v1

@@ 1 Introduction
The hierarchical Bayesian formulations for population inference usually treat either (i) a set of statistically independent, individually resolved events, or (ii) a stochastic signal that can be related to the underlying source population. However, they do not accommodate situations where a single population contributes simultaneously to a resolvable population and to an unresolved confusion background - precisely the case for LISA GBs, and potentially for EMRIs. A formalism that includes both resolved events and stochastic backgrounds was proposed in the LVK context [73], but it does not account for the correlations that naturally emerge in a LISA-like Global Fit, particularly the fact that the number of resolvable events is not fixed and that resolving more or fewer sources directly impacts the inferred foreground. As LISA data-analysis tools are being developed, it is becoming essential to establish a unified formalism capable of treating such populations, both to guide methodological development and to define the data products that should be provided by the collaboration. This issue is also relevant for third-generation ground-based detectors, which will come online in the next decade, such as the Einstein Telescope [74] and Cosmic Explorer [75], which are expected to detect an astrophysical stochastic background [76, 77, 78, 79] in addition to individual events, and for pulsar timing arrays, which are expected to resolve individual sources in the near future [80, 81, 82, 83], particularly once the Square Kilometre Array [84] becomes operational.
@@ 2 Current hierarchical Bayesian formalism
$p(\{d_{c}\}_{n_{\rm det}}|\Lambda)\propto e^{-\xi(\Lambda)R}\prod_{i=1}^{n_{\rm det}}\int p(d_{c,i}|\theta)N(\theta_{i}|\Lambda)\ {\rm d}\theta.$ (2.4)
@@
$p(d|\Lambda)\propto e^{-\xi(\Lambda)\,R}\prod_{i=1}^{n_{\rm det}}\int p(d_{c,i}|\theta,{\rm obs})\,N(\theta_{i}|\Lambda)\,{\rm d}\theta\times\int p(d|\Sigma)\,p(\Sigma|\Lambda)\,{\rm d}\Sigma\,.$ (2.8)
In the above, the second term dependence on the full data is only an approximation, as it includes also all of the resolved-source containing segments, $\{d_{c,i}\}$. It is a good approximation for current ground-based detectors because data chunks containing detectable events are much rarer than those without. We can therefore assume that these rare events do not affect the reconstruction of the stochastic noise, and conversely, that the noise can be considered known when these events are observed, and does not need to be constrained jointly. In App. A, we discuss the relation between our formalism and Eq. (2.8).
@@ 5 Conclusions
In this work, we introduced an approach that coherently incorporates the correlation between the resolvable population and the background, as well as their joint relation to the astrophysical population, without requiring a post-processed catalogue of sources extracted from the Global Fit. In the Appendix, we also show how this general framework can approach the current LVK formalism, and discuss differences with the formalism proposed in [73] to combine background and individual events.
@@ Appendix A LVK limit
The underlying assumption of the population analysis as used in the LVK is that the data has been divided into segments, and that a very large number of segments, $K\gg 1$, have been observed. Denoting the expected number of events in the Universe that would occur during the observation time by $N$, the data in segment $i$ by $d_{c,i}$, and introducing flags $\{f_{i}\}$ that indicate the segments that contain ($f_{i}=1$) or do not contain ($f_{i}=0$) events, the likelihood can be written as
$\prod_{i=1}^{K}\left[f_{i}p(d_{c,i}|\theta_{i})+(1-f_{i})p_{n}(d_{c,i})\right]$ (A.1)
@@
$\prod_{i=1}^{K}\left[\frac{N}{K}p(d_{c,i}|\theta_{i})+\left(1-\frac{N}{K}\right)p_{n}(d_{c,i})\right].$ (A.2)
To obtain the posterior on the parameters of the individual sources and population parameters we now introduce priors on the event parameters. In the LVK analysis we additionally assume that only segments with data satisfying our selection criterion contain information about the population. We additionally assume that in those segments $p(d_{c,i}|\theta_{i})\gg p_{n}(d_{c,i})$. Marginalising over the data and parameters in the other segments gives us
$\left[\prod_{i=1}^{N_{\rm obs}}\frac{N}{K}p(d_{c,i}|\theta_{i})\right]\left[\frac{N}{K}p_{\rm ndet}(\Lambda)+\left(1-\frac{N}{K}\right)\right]^{K-N_{\rm obs}}$
$=\left[\prod_{i=1}^{N_{\rm obs}}\frac{N}{K}p(d_{c,i}|\theta_{i})\right]\exp(Np_{\rm ndet}(\Lambda)-N),$ (A.3)
where the last line follows from taking the limit as $K\rightarrow\infty$. We see that we recover the usual expression as expected.
In our formalism we have two separate processes, generating resolved and background events, and so we introduce a second flag, $b_{i}$, which is $1$ for segments that include unresolved events, and 0 for segments that do not. The first flag, $f_{i}$, will indicate the presence or absence of resolved events. In the LVK limit, it is assumed that events are rare and so segments cannot contain both resolved and background events. However, we do not impose this a priori, but the cross-terms it introduces will disappear when we take the limit $K\rightarrow\infty$.
$\prod_{i=1}^{K}\left[f_{i}p(d_{c,i}|\theta_{i})+(1-f_{i})g_{i}p(d_{c,i}|\theta_{i})+(1-f_{i})(1-g_{i})p_{n}(d_{c,i})\right]$ (A.4)
The priors on the two sets of flags are binomial distributions with $p=N_{1}/K$ and $p=N_{2}/K$ respectively. The priors on the parameter values are $p_{1}(\theta|\Lambda)$ and $p_{2}(\theta|\Lambda)$. Marginalising over both the flags and parameters of individual sources we obtain
$\prod_{i=1}^{K}\left[\frac{N_{1}}{K}p_{1}(d_{c,i}|\Lambda)+\left(1-\frac{N_{1}}{K}\right)\frac{N_{2}}{K}p_{2}(d_{c,i}|\Lambda)+\left(1-\frac{N_{1}}{K}\right)\left(1-\frac{N_{2}}{K}\right)p_{n}(d_{c,i})\right]$ (A.5)
@@
$p_{1,2}(d_{c,i}|\Lambda)=\int p_{1,2}(d_{c,i}|\theta_{i})p_{1,2}(\theta|\Lambda).$ (A.6)
Note, that in a typical situation the likelihood does not differ between resolved and unresolved sources, so $p_{1}(d_{c,i}|\theta_{i})\equiv p_{2}(d_{c,i}|\theta_{i})$. In the above we allow for the likelihoods to be different because it is required to achieve an LVK-like separation in our formalism, which we discuss further below.
@@
$\left(1-\frac{N_{1}}{K}\right)^{K-N_{\rm obs}}\prod_{j=1}^{K-N_{\rm obs}}\left[\frac{N_{2}}{K}p_{2}(d_{c,j}|\Lambda)+\left(1-\frac{N_{2}}{K}\right)p_{n}(d_{c,j})\right].$ (A.7)
@@
The form of the model used in our formalism is inherently probabilistic. All data segments are sometimes used for inference of resolved sources and sometimes not. It is therefore not possible to directly recover the LVK limit unless we impose additional assumptions (e.g., selection of data) or assume that the likelihood takes a somewhat contrived form (this will be done in the final subsection as an illustration). But we can attempt to apply some of the simplifications that are used to obtain the LVK population inference formalism to our framework to obtain results that are similar to the LVK.
Working with Eq. (A.4) and obtaining the posterior on a given flag, $f_{i}$, we can see that the ratio of the probability that $f_{i}=1$ to the probability that $f_{i}=0$ is just the Bayes factor for the resolved source model versus the background plus instrumental noise model. This Bayes factor depends only on the data and the population parameters $\Lambda$ and not on the parameters of the potential source in the segment. For fixed population parameters, it can be used for data selection, similar to the data selection used in the LVK analysis. If we impose some threshold on this Bayes factor then we separate data segments into those above threshold and those below threshold. If we assume additionally that this threshold is high enough that the first term in the brackets dominates for those data segments, these terms contribute
$\prod_{i=1}^{N_{\rm obs}}\frac{N_{1}}{K}p_{1}(d_{c,i}|\Lambda).$ (A.8)
This can be simplified by noting
$p_{1,2}(d_{c}|\Lambda)$ $=\frac{N}{N_{1,2}}\tilde{p}_{1,2}(d_{c}|\Lambda)$
$\mbox{where}\qquad\tilde{p}_{1}(d_{c}|\Lambda)$ $=\int p(d_{c}|\theta)p(\theta|\Lambda)\alpha(\theta,\Lambda){\rm d}\theta$
$\tilde{p}_{2}(d_{c}|\Lambda)$ $=\int p(d_{c}|\theta)p(\theta|\Lambda)(1-\alpha(\theta,\Lambda)){\rm d}\theta$ (A.9)
and so the resolved source term becomes
$\prod_{i=1}^{N_{\rm obs}}\frac{N}{K}\tilde{p}_{1}(d_{c,i}|\Lambda).$ (A.10)
@@
$\int\frac{N}{K}(\tilde{p}_{1}(\mathbf{d}|\Lambda)+\tilde{p}_{2}(\mathbf{d}|\Lambda))+\left(1-\frac{N}{K}\right)p_{n}(\mathbf{d}){\rm d}\mathbf{d}$ (A.11)
where we have dropped terms that are quadratic in $N/K$ and the integral is over data segments that are “not detected”, i.e., below threshold. If we assume that the threshold is high enough that noise cannot generate data in the detection-region the integral of $p_{n}(d_{c})$ over the non-detection region is equal to unity and this term becomes
$\frac{N}{K}\left(1-p_{\rm det}(\Lambda)\right)+\left(1-\frac{N}{K}\right)=1-\frac{N}{K}p_{\rm det}(\Lambda)$ (A.12)
where
$p_{\rm det}(\Lambda)=\int_{d_{c}\in\mbox{detected}}p(d_{c}|\theta)p(\theta|\Lambda){\rm d}d_{c}.$ (A.13)
There are $K-N_{\rm obs}$ of these terms, and so we have
$\left(1-\frac{N}{K}p_{\rm det}(\Lambda)\right)^{K-N_{\rm obs}}\rightarrow\exp(-Np_{\rm det}(\Lambda))\qquad\mbox{as}\qquad K\rightarrow\infty.$ (A.14)
We see that we recover the standard LVK expression, with the exception that the resolved sources contribute $\tilde{p}_{1}(d_{c}|\Lambda)$, which depends on $\alpha(\theta,\Lambda)$. However, with the assumption that the segments above threshold clearly contain resolved sources, alpha must be approximately equal to 1 for all the parameter values generating data in this region.
In the above we have assumed that segments contain not only at most one resolved source but also at most one background source. This is consistent with the assumptions used by the LVK, but we would usually expect backgrounds to be always present. This assumption is not required, however, and can be eliminated by replacing Eq. (A.5) with
@@
$p_{m}(d_{c}|\Lambda)$ $=\int p_{n}\left(d_{c}-\sum_{i=1}^{m}h(\theta_{i})\right)\prod_{i=1}^{m}\frac{p(\theta_{i}|\Lambda)(1-\alpha(\theta_{i},\Lambda))}{N_{2}/N}{\rm d}\theta_{i}.$ (A.16)
Proceeding as before and again assuming that the probability that the background process generates data in the detected region is vanishingly small, so that the integral of $q(d_{c}|\Lambda)$ over that region is $1$, we obtain
$\frac{N}{K}\left(\frac{N_{1}}{N}-p_{1,{\rm det}}(\Lambda)\right)+\left(1-\frac{N}{K}\right)=1-\frac{N}{K}p_{1,{\rm det}}(\Lambda)$ (A.17)
@@
$p_{1,{\rm det}}(\Lambda)=\int_{d_{c}\in\mbox{detected}}p(d_{c}|\theta)p(\theta|\Lambda)\alpha(\theta,\Lambda){\rm d}d_{c}.$ (A.18)
Under the assumptions we have made, we can assume that $\alpha\approx 1$ in the region of parameter space generating data in the detectable region and so $p_{1,{\rm det}}(\Lambda)\approx p_{{\rm det}}(\Lambda)$ and so we once again obtain the standard result.
@@
$\prod_{i=1}^{K-N_{\rm obs}}\left[\frac{N}{K}\tilde{p}_{1}(d_{c,i}|\Lambda)+\left(1-\frac{N_{1}}{K}\right)q(d_{c,i}|\Lambda)\right].$ (A.19)
@@
To do things correctly we need to explicitly keep all these terms, but sampling all of them would be expensive. One approximation would be to assume that for all of these terms only the $q(d_{c}|\Lambda)$ part is important. This is not true, however, as the transition from resolved to background sources should not be abrupt (although see next section). If we did do this the prefactors $(1-N_{1}/K)^{K-N_{\rm obs}}$ would contribute a term $\exp(-N_{1})$. However, this is not consistent since we are keeping some terms proportional to $N_{1}/K$ and not others. Alternatively, the resolved source term could be marginalised over data, so that these segments do not have to be fully analysed with a resolved source model, but we retain the data dependence in the second term. This is computationally simpler, but also somewhat inconsistent. A final approach would be to approximate this product of a large number of terms as a Gaussian, appealing to the central limit theorem. The contribution to the Gaussian variance would then come from both resolved and background events and hence depend on the total rate $N$. No matter how this term is treated we note that it does not give the $\exp(-\xi(\Lambda)R)$ term that appears in Eq. (2.8). However, we believe this is because that term should not be there. In the standard derivation of the LVK population likelihood, this term arises because of the marginalisation over the data segments that are not being used for the resolved source analysis. Once we add the background term $p(d|\Sigma)$ we are reintroducing that data and so it has not been marginalised out, therefore this term should no longer appear.
A.1 Recovering automatic data selection
In the previous section we showed how results similar to the LVK can be obtained by marginalising out some of the data segments. To do this we had to impose data selection (based on the Bayes factor) on top of the split between resolved and unresolved events. We show in the following how to build a contrived model in which this additional condition is not needed, by ensuring that the Bayes factor sharply transitions from close to zero to a very large value at a certain data threshold. Under these circumstances, segments below that threshold will effectively only use the background likelihood and those above will only use the resolved source likelihood, and we obtain an LVK-like expression in which there is no ambiguity between which segments contain resolved sources and which do not, and this disticntion corresponds to data selection in the LVK style.
This data selection will happen automatically in our formalism if and only if there is a region of data space where $N_{1}p_{1}(d_{c}|\Lambda)\gg N_{2}p_{2}(d_{c}|\Lambda)$ and $N_{1}p_{1}(d_{c}|\Lambda)\gg Kp_{n}(d_{c})$ and the converse is true in all other regions of data space. Assuming this is the case then any segments that lie in this region will essentially contribute only through the first term in the sum, while all other segments will only contribute through the second term in the sum. If such a likelihood exists then we can induce an LVK-like selection based only on properties of the data. Constructing such a model is hard, because the resolved and background populations share sources with the same parameters, but we need to ensure that those events give different likelihoods in the two cases. A contrived way to do this is to suppose that there are two different likelihoods for a given set of parameters, which we call the resolved likelihood, $p_{\rm res}(d_{c}|\theta)$ and the background likelihood, $p_{\rm b}(d_{c}|\theta)$. An event with parameters $\theta$ has probability $\alpha(\theta,\Lambda)$ of generating data from the resolved likelihood, otherwise it generates data from the background likelihood. Assuming that the supports of $p_{\rm res}(d_{c}|\theta)$ and $p_{\rm b}(d_{c}|\theta)$ are distinct, we achieve the desired separation. Obviously this is highly contrived, but it is the only way to get a strict separation between the two populations in the manner of the standard LVK analysis in our formalism.
Under these assumptions, the likelihood picks out the $N_{\rm obs}$ segments with data lying in the resolved region

### v2

@@ 1 Introduction
The hierarchical Bayesian formulations for population inference usually treat either (i) a set of statistically independent, individually resolved events, or (ii) a stochastic signal that can be related to the underlying source population. However, they do not accommodate situations where a single population contributes simultaneously to a resolvable population and to an unresolved confusion background - precisely the case for LISA GBs, and potentially for EMRIs. A formalism that includes both resolved events and stochastic backgrounds was proposed in the LVK context [73], but it does not treat resolved and unresolved sources in a fully consistent way. As a consequence, it does not account for the correlations that naturally emerge in a LISA-like Global Fit, particularly the fact that the number of resolvable events is not fixed and that resolving more or fewer sources directly impacts the inferred background. As LISA data-analysis tools are being developed, it is becoming essential to establish a unified formalism capable of treating such populations, both to guide methodological development and to define the data products that should be provided by the collaboration. This issue is also relevant for third-generation ground-based detectors, which will come online in the next decade, such as the Einstein Telescope [74] and Cosmic Explorer [75], which are expected to detect an astrophysical stochastic background [76, 77, 78, 79] in addition to individual events, and for pulsar timing arrays, which are expected to resolve individual sources in the near future [80, 81, 82, 83], particularly once the Square Kilometre Array [84] becomes operational.
@@ 2 Current hierarchical Bayesian formalism
$p(\{d_{c}\}_{n_{\rm det}}|\Lambda)\propto e^{-\xi(\Lambda)N}\prod_{i=1}^{n_{\rm det}}\int p(d_{c,i}|\theta)N(\theta_{i}|\Lambda)\ {\rm d}\theta.$ (2.4)
@@
$p(d|\Lambda)\propto e^{-\xi(\Lambda)N}\prod_{i=1}^{n_{\rm det}}\int p(d_{c,i}|\theta,{\rm obs})\,N(\theta_{i}|\Lambda)\,{\rm d}\theta\times\int p(d|\Sigma)\,p(\Sigma|\Lambda)\,{\rm d}\Sigma\,.$ (2.8)
In the above, the second term dependence on the full data is only an approximation, as it includes also all of the resolved-source containing segments, $\{d_{c}\}_{n_{\rm det}}$. This is acceptable for current ground-based detectors because data chunks containing detectable events are much rarer than those without. We can therefore assume that these rare events do not affect the reconstruction of the stochastic noise, and conversely, that the noise can be considered known when these events are observed, and does not need to be constrained jointly. However, as we will show in App. A, this expression does not fully incorporate the information of resolved and unresolved sources self-consistently. It is a good approximation for the current situation, where we can only place upper bounds on the background.
@@ 5 Conclusions
In this work, we introduced an approach that coherently incorporates the correlation between the resolvable population and the background, as well as their joint relation to the astrophysical population, without requiring a post-processed catalogue of sources extracted from the Global Fit. In the Appendix, we also show how this general framework can approach the current LVK formalism, and discuss differences with the formalism proposed in [73] to combine background and individual events. We outline how this formalism should evolve as the number of detected events and the overlap between them grow, particularly in anticipation of third-generation detectors such as the Einstein Telescope or Cosmic Explorer.
@@ Appendix A LVK limit
The underlying assumption of the population analysis as used in the LVK is that the data has been divided into segments, and that a very large number of segments, $K\gg 1$, have been observed. We apply our treatment of probabilistic inclusion of sources to each segment, first considering only resolvable-like sources. Denoting the expected number of events in the Universe that would occur during the observation time by $N$, the data in segment $i$ by $d_{c,i}$, and introducing flags $\{f_{i}\}$ that indicate the segments that contain ($f_{i}=1$) or do not contain ($f_{i}=0$) events, the joint likelihood can be written as
$p(d,\{\theta\}_{K}|\Lambda)\propto\prod_{i=1}^{K}\left[f_{i}p(d_{c,i}|\theta_{i})p(\theta_{i}|\Lambda)+(1-f_{i})p_{n}(d_{c,i})\right]$ (A.1)
@@
$p(d,\{\theta\}_{K}|\Lambda)\propto\prod_{i=1}^{K}\left[\frac{N}{K}p(d_{c,i}|\theta_{i})p(\theta_{i}|\Lambda)+\left(1-\frac{N}{K}\right)p_{n}(d_{c,i})\right].$ (A.2)
The form of the model used in our formalism is inherently probabilistic. All data segments are sometimes used for inference of resolved sources and sometimes not. The standard LVK approach on the other hand discards data segments that are thought to contain no events based on some criterion, i.e.,  data is selected. Working with Eq. (A.1) and obtaining the posterior on a given flag, $f_{i}$, we can see that the ratio of the probability that $f_{i}=1$ to the probability that $f_{i}=0$ is just the Bayes factor for the resolved source model versus instrumental noise model in a given segment. This Bayes factor depends on the data (i.e., on the parameters that generated the event and on the noise realisation, see [93]), but also on the population parameters $\Lambda$. The LVK approach uses a fixed population model for computing the detection statistic, but in principle selection on data should be carried out jointly with the hierarchical inference. Assuming the threshold is chosen high enough that the population model has relatively little impact on which segments are selected, then we can separate data segments into those that are above this threshold and those that are below the threshold. We additionally assume that in the segments above threshold $p(d_{c,i}|\theta_{i})\gg p_{n}(d_{c,i})$. Marginalising over the data and parameters in the other segments, and assuming that the threshold has been chosen so that the noise generating likelihood has no support in the “detected” data region, gives us
$p(\{d_{c}\}_{n_{\rm det}}|\Lambda)$ $\propto\left[\prod_{i=1}^{n_{\rm det}}\frac{N}{K}\int p(d_{c,i}|\theta)p(\theta|\Lambda){\rm d}\theta\right]\left[\frac{N}{K}p_{\rm ndet}(\Lambda)+\left(1-\frac{N}{K}\right)\right]^{K-n_{\rm det}}$
$\propto\left[\prod_{i=1}^{n_{\rm det}}\frac{N}{K}\int p(d_{c,i}|\theta)p(\theta_{i}|\Lambda){\rm d}\theta\right]e^{Np_{\rm ndet}(\Lambda)-N},$ (A.3)
where the last line follows from taking the limit as $K\rightarrow\infty$, and we introduced
$p_{\rm ndet}(\Lambda)$ $=\int_{d_{c}\in{\rm ndet}}p(d_{c}|\theta)p(\theta|\Lambda)\ {\rm d}d_{c}{\rm d}\theta$ (A.4)
$=1-\xi(\Lambda)$
Finally, we get
$p(\{d_{c}\}_{n_{\rm det}}|\Lambda)\propto e^{-N\xi(\Lambda)}\prod_{i=1}^{n_{\rm det}}\int p(d_{c,i}|\theta)N(\theta_{i}|\Lambda){\rm d}\theta\\$ (A.5)
We see that we recover the usual expression as expected.
We now include a second process that generates background-like events, with also at most one such event per segment. We discuss afterwards the more realistic case of having multiple background-contributing events per segment. We allow a priori for having both resolved and background events in the same segment, but the cross-terms this introduces will disappear when we take the limit $K\rightarrow\infty$. Following our previous notation, we note the population priors on resolved and background events $p_{1}(\theta|\Lambda)$ and $p_{2}(\theta|\Lambda)$, respectively. They are related to the global population prior through the resolvability function $\alpha(\theta,\Lambda)$, as in Eqs. (3.12) and (3.13). We introduce a second flag, $b_{i}$, which is $1$ for segments that include unresolved events, and 0 for segments that do not. The first flag, $f_{i}$, will indicate the presence or absence of resolved events. The joint likelihood now reads
$p(d,\{\theta\}^{1}_{K}|\Lambda),\{\theta\}^{2}_{K}|\Lambda)\propto\prod_{i=1}^{K}\left[f_{i}p_{1}(d_{c,i}|\Lambda)+(1-f_{i})g_{i}p(d_{c,i}|\Lambda)+(1-f_{i})(1-g_{i})p_{n}(d_{c,i})\right],$ (A.6)
@@
$p_{1,2}(d_{c}|\Lambda)=\int p_{1,2}(d_{c}|\theta_{i},\Lambda)p_{1,2}(\theta|\Lambda).$ (A.7)
Note, that in a typical situation the likelihood does not differ between resolved and unresolved sources, so $p_{1}(d_{c,i}|\theta_{i})\equiv p_{2}(d_{c,i}|\theta_{i})$. In the above we allow for the likelihoods to be different because this allows for greater generality. We include a $\Lambda$ dependency in the likelihoods, as, in all generality, the total noise model used there depends on $\Lambda$ through the background.
The priors on the two sets of flags are binomial distributions with $p=N_{1}/K$ and $p=N_{2}/K$ respectively. Marginalising over both the flags and parameters of individual sources we obtain
$p(d,\{\theta\}^{1}_{K}|\Lambda),\{\theta\}^{2}_{K}|\Lambda)$
$\propto\prod_{i=1}^{K}\left[\frac{N_{1}}{K}p_{1}(d_{c,i}|\Lambda)+\left(1-\frac{N_{1}}{K}\right)\frac{N_{2}}{K}p_{2}(d_{c,i}|\Lambda)+\left(1-\frac{N_{1}}{K}\right)\left(1-\frac{N_{2}}{K}\right)p_{n}(d_{c,i})\right].$ (A.8)
@@
$\left(1-\frac{N_{1}}{K}\right)^{K-n_{\rm det}}\prod_{j=1}^{K-n_{\rm det}}\left[\frac{N_{2}}{K}p_{2}(d_{c,j}|\Lambda)+\left(1-\frac{N_{2}}{K}\right)p_{n}(d_{c,j})\right].$ (A.9)
@@
The ratio of the probability that $f_{i}=1$ to the probability that $f_{i}=0$ is now the Bayes factor for the resolved source model versus the background plus instrumental noise model. We threshold as before on this Bayes factor, and separate data segments into those above threshold and those below threshold. If we assume additionally that this threshold is high enough that the resolved source term dominates for those data segments, these terms contribute
$\prod_{i=1}^{n_{\rm det}}\frac{N_{1}}{K}p_{1}(d_{c,i}|\Lambda).$ (A.10)
This can be simplified by defining
$\tilde{p}_{1,2}(d_{c}|\Lambda)$ $=\frac{N_{1,2}}{N}p_{1,2}(d_{c}|\Lambda),$ (A.11)
which is just the numerator part of Eqs. (3.12) and (3.13). The resolved source term becomes
$\prod_{i=1}^{n_{\rm det}}\frac{N}{K}\tilde{p}_{1}(d_{c,i}|\Lambda).$ (A.12)
@@
$\int_{d_{c}\in{\rm ndet}}\frac{N}{K}\left(\tilde{p}_{1}(d_{c}|\Lambda)+\tilde{p}_{2}(d_{c}|\Lambda)\right)+\left(1-\frac{N}{K}\right)p_{n}({d_{c}})\ {\rm d}{d_{c}}$ (A.13)
where we have dropped terms that are quadratic in $N/K$. If we assume that the threshold is high enough that noise cannot generate data in the detection-region the integral of $p_{n}(d_{c})$ over the non-detection region is equal to unity and this term becomes
$\frac{N}{K}\left(1-\xi(\Lambda)\right)+\left(1-\frac{N}{K}\right)=1-\frac{N}{K}\xi(\Lambda),$ (A.14)
which follows because the integral of $\tilde{p}_{1}(d_{c}|\Lambda)+\tilde{p}_{2}(d_{c}|\Lambda)$ yields the total fraction of non detected events $1-\xi(\Lambda)$). There are $K-N_{\rm obs}$ of these terms, and so we have
$\left(1-\frac{N}{K}\xi(\Lambda)\right)^{K-n_{\rm det}}\rightarrow e^{-N\xi(\Lambda)}\qquad\mbox{as}\qquad K\rightarrow\infty.$ (A.15)
Putting the pieces together, we get
$p(\{d_{c}\}_{n_{\rm det}}|\Lambda)\propto e^{-N\xi(\Lambda)}\prod_{i=1}^{n_{\rm det}}\int p_{1}(d_{c,i}|\Lambda)N_{1}(\theta|\Lambda).$ (A.16)
We see that we recover the standard LVK expression, with the exception that the resolved sources contribute $p_{1}(\theta|\Lambda)$, which depends on $\alpha(\theta,\Lambda)$. However, with the assumption that the segments above threshold clearly contain resolved sources, alpha must be approximately equal to 1 for all the parameter values generating data in this region.
In the above we have assumed that segments contain not only at most one resolved source but also at most one background source. This assumption is not required, however, and can be eliminated by replacing Eq. (A.8) with
@@
$p_{m}(d_{c}|\Lambda)$ $=\int p_{n}\left(d_{c}-\sum_{i=1}^{m}h(\theta_{i})\right)\prod_{i=1}^{m}\frac{p(\theta_{i}|\Lambda)(1-\alpha(\theta_{i},\Lambda))}{N_{2}/N}{\rm d}\theta_{i},$ (A.18)
with $h(\theta)$ being the GW signal of a source. Marginalising over non-detectable data as before, and again assuming that the probability that the background process generates data in the detected region is vanishingly small, so that the integral of $q(d_{c}|\Lambda)$ over that region is $1$, we obtain instead of Eq. (A.19)
$\frac{N}{K}\left(\frac{N_{1}}{N}-\xi_{1}(\Lambda)\right)+\left(1-\frac{N}{K}\right)=1-\frac{N}{K}\xi_{1}(\Lambda)$ (A.19)
@@
$\xi_{1}(\Lambda)=\int_{d_{c}\in{\rm det}}p(d_{c}|\theta)p(\theta|\Lambda)\alpha(\theta,\Lambda){\rm d}\ d_{c}.$ (A.20)
Under the assumptions we have made, we can assume that $\alpha\approx 1$ in the region of parameter space generating data in the detectable region and so $\xi_{1}(\Lambda)\approx\xi(\Lambda)$ and so we once again obtain the standard result.
@@
$\prod_{i=1}^{K-n_{\rm det}}\left[\frac{N}{K}\tilde{p}_{1}(d_{c,i}|\Lambda)+\left(1-\frac{N_{1}}{K}\right)q(d_{c,i}|\Lambda)\right].$ (A.21)
@@
To do things correctly we need to explicitly keep all these terms, but sampling all of them would be expensive. One approximation would be to assume that for all of these terms only the $q(d_{c}|\Lambda)$ part is important. This is not true, however, as the transition from resolved to background sources should not be abrupt (although see next section). If we did do this the prefactors $(1-N_{1}/K)^{K-n_{\rm det}}$ would contribute a term $\exp(-N_{1})$. However, this is not consistent since we are keeping some terms proportional to $N_{1}/K$ and not others. Alternatively, the resolved source term can be averaged over data, so that these segments do not have to be fully analysed with a resolved source model, but we retain the data dependence in the second term. To do this we write the above expression as
$\left(1-\frac{N_{1}}{K}\right)^{K-n_{\rm det}}\prod_{i=1}^{K-n_{\rm det}}q({d}_{c,i}|\Lambda)\prod_{j=1}^{K-n_{\rm det}}\left[1+\frac{N}{(K-N_{1})}\frac{\tilde{p}_{1}({d}_{c,i}|\Lambda)}{q({d}_{c,i}|\Lambda)}\right]$